# Electric-Field Driven Nuclear Dynamics of Liquids and Solids from a Multi-Valued Machine-Learned Dipolar Model

### v1

@@ V.2 LiNbO3
The LiNbO3 dataset for training the MACE MLIP and dipole models was obtained from $3$ ab initio NVT trajectories of 20 ps at 500 K, 1000 K, 1500 K (SVR thermostat, $\tau=100$ fs, $dt=2$ fs) with a simulation box containing $30$ atoms and lattice vectors $a,b,c=5.14,5.14,13.82$  and angles $\alpha,\beta,\gamma=90^{\circ},90^{\circ},120^{\circ}$. We used the FPS algorithm to select $5000$ structures. For these structures, we computed the polarization using the Berry-phase formalism implemented in FHI-aims [69]. The values of the dipoles have been branch-matched as described in Supplementary Section S9.
The MLIP consists of a MACE model with a cutoff radius of $6$, $2$ layers, $l=2$ as the highest spherical harmonics, a maximum of $l=0$ for each message, correlation order $3$ at each layer. The parameters for the MACE dipole model were the same as the ones used for water, but adjusting for the oxidation numbers of LiNbO3. The final errors on energies, forces and dipoles for these models are reported in Table 3.
The simulations regarding the ferroelectric to paraelectric phase transition in Section III.2 were obtained by using a $4\!\times\!4\!\times\!$ 4 × 4 supercell with $640$ atoms, lattice vectors of length $21.94$  and angles between them of $55.83^{\circ}$. For each temperature we simulated $10$ independent NVT runs (Langevin thermostat, $\tau=100$ fs, $dt=1$ fs). We used the trained dipole model to evaluate the dipole for each step.
The simulation with a time-dependent laser pulse, whose results are reported in Fig. 3c, were run with the same $4\!\times\!4\!\times\!$ 4 × 4 supercell as used above, and were run in the NVE ensemble, starting from a thermalized trajectory. We used a time-step of 0.1fs to ensure stability of the integrator with a time-dependent Hamiltonian. For each pulse considered, we ran 16 independent trajectories starting from randomly sampled geometries equilibrated at 300 K. The average and standard deviation at varying time shown in Fig. 3c correspond to the statistical analysis over these independent trajectories.
@@ Acknowledgments
E.S. thanks Federico Ernesto Mocchetti for the mathematical background regarding the topological aspects of the multi-valued ML models and Shubham Sharma for discussions and help with training MACE models. We thank George Trenins for a critical read of the paper draft and insightful discussions. We thank Vasily Artemov for sending us a table with the experimental datapoints of the dielectric permittivity of water. We also thank Michael Fechner for insightful discussions about phonon dynamics in LiNbO3.

### v2

@@ IV.4 LiNbO3 Simulation Details
The LiNbO3 dataset for training the MACE MLIP and dipole models was obtained from $3$ ab initio NVT trajectories of 20 ps at 500 K, 1000 K, 1500 K (SVR thermostat, $\tau=100$ fs, $dt=2$ fs) with a simulation box containing $30$ atoms and lattice vectors $a,b,c=5.14,5.14,13.82$ Å  and angles $\alpha,\beta,\gamma=90^{\circ},90^{\circ},120^{\circ}$. We used the FPS algorithm to select $5000$ structures. For these structures, we computed the polarization using the Berry phase formalism implemented in FHI-aims [74]. The values of the dipoles have been branch-matched as described in Supplementary Section S10.
The MLIP consists of a MACE model with a cutoff radius of $6$ Å, $2$ layers, $\ell=2$ as the highest spherical harmonics, a maximum of $L=0$ for each message, correlation order $3$ at each layer. The parameters for the MACE dipole model were the same as the ones used for water, but adjusting for the oxidation numbers of LiNbO3. The final errors on energies, forces and dipoles for these models are reported in Table 3.
The simulations regarding the ferroelectric to paraelectric phase transition in Section II.3 were obtained by using a $4\!\times\!4\!\times\!4$ supercell with $640$ atoms, lattice vectors of length $21.94$ Å  and angles between them of $55.83^{\circ}$. For each temperature we simulated $10$ independent NVT runs (Langevin thermostat, $\tau=100$ fs, $dt=1$ fs). We used the trained dipole model to evaluate the dipole for each step.
The simulation with a time-dependent laser pulse, whose results are reported in Fig. 3c, were run with the same $4\!\times\!4\!\times\!4$ supercell as used above, and were run in the NVE ensemble, starting from a thermalized trajectory. We used a time-step of 0.1fs to ensure stability of the integrator with a time-dependent Hamiltonian. For each pulse considered, we ran 16 independent trajectories starting from randomly sampled geometries equilibrated at 300 K. The average and standard deviation at varying time shown in Fig. 3c correspond to the statistical analysis over these independent trajectories.
@@ Acknowledgments
E.S. thanks Federico Ernesto Mocchetti for the mathematical background regarding the topological aspects of the multi-valued ML models and Shubham Sharma for discussions and help with training MACE models. We thank George Trenins for a critical read of the paper draft and insightful discussions. We thank Vasily Artemov for sending us a table with the experimental datapoints of the dielectric permittivity of water. We also thank Michael Fechner for insightful discussions about phonon dynamics in LiNbO3. M.R. acknowledges funding by the European Union (ERC, QUADYMM, 101169761).
@@
Competing Interests
The authors declare no competing interests.